OpenStudy (anonymous):

An original rectangle has a length of 8 and a width of 24. What happens to the area if the new width is one half the original width? Verify your answer.

OpenStudy (anonymous):

the area of the rectangle becomes half.

OpenStudy (anonymous):

initial area = 8*24= 192 sq units if width is 12 unit area = 8*12 = 96 sq units so area is half

OpenStudy (amistre64):

Area = length * width\[\] If width is reduced by half we get:\[\] Area = length * width/2 which is that same as A = (1w)/2\[\] its half of what it was